Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
nardecchia, gherardini - 22:46 Thursday 30 May 2024 (64386) Print this report
NI CO2 laser noise injections

CO2 noise injection survey has been performed this afternoon to investigate the possibility of exciting some mechanical modes of the CP using the CO2 laser piezo with the current DAS injected power (~ 4 W ) . 

With the ITF stably locked in LN3, we injected a single line on the NI CO2 laser at a fixed frequency for 5 minutes (amplitude applied on the piezo=4V), then we moved to the next one covering the frequency range from 100Hz to 3kHz.

There wasn't enough time to inject noise at all planned frequencies (we reached ~1 kHz).
The noise lines clearly appeared in hrec but nothing unusual was seen. The comparison between the measured hrec values and those predicted by the model will be done in the following days.

In the attached file, frequency, start time and gps of noise injections are listed.

External Note: I needed to re-adjust the WI Chiller setpoint to recover the standard laser condition (set point=19.03 degrees at 20.41 UTC) 


Non-image files attached to this report
Comments to this report:
nardecchia, lorenzini - 18:39 Friday 31 May 2024 (64393) Print this report

In this entry, the results of the CO2 noise injections are reported. 

The injected lines in hrec are shown in figures 1, 2 and 3.  The measured hrec values and the predicted ones are listed in the attached pdf file.

The comparison between the measured and expected hrec values is shown in fig.4. 

The two curves differ roughtly by a factor ~3 as we already measured in the past CO2 noise injections (63287).

Note that the line at 1030 Hz is invisible. 


Images attached to this comment
Non-image files attached to this comment
Search Help